OpenStudy (anonymous):

2/3(x − 3) = 1/6(8x + 31) + 3 Solve Thanks in advance!

OpenStudy (rajee_sam):

\[\frac{ 2 }{ 3 (x-3) } = \frac{ 1 }{ 6(8x + 31) +3 }\] is this the question?

OpenStudy (rajee_sam):

@Oglon3r

OpenStudy (anonymous):

\[\frac{ 2 }{ 3 } (x-3)=\frac{ 1 }{ 6 }(8x+31)+3\]

OpenStudy (anonymous):

that one thank you for the reply

OpenStudy (rajee_sam):

OK

OpenStudy (anonymous):

x=-61/4

OpenStudy (anonymous):

oh dang I got a completely different answer... mind if I ask how that is the answer?

OpenStudy (rajee_sam):

\[\frac{ 4 }{ 6 } (x-3) = \frac{ 1 }{ 6 } (8x +1) + \frac{ 18 }{ 6 }\] =\[4(x-3) = (8x + 1) + 18\]\[4x - 12 = 8x + 1 + 18 \]\[-12-19 = 4x\] \[4x = -31 \]\[x = \frac{ -31 }{ 4 }\]

OpenStudy (anonymous):

now it is wrong at the first step itself

OpenStudy (rajee_sam):

which is??

OpenStudy (anonymous):

opps sorry